An electric heater is used to supply 54.3 J of energy to a 15.0-g graphite block with an initial temperature of 20.0 ºC. Determine the final temperature of the graphite. (Specific heat of graphite = 0.72 J/g ºC.)
1
Expert's answer
2020-05-06T13:58:11-0400
"Q = cm\Delta T"

"54.3 = 0.72*15.0*\Delta T"

"\Delta T = {54.3 \over 0.72*15.0} = 5.028 ^oC \approx 5.0 ^oC"

"T_2 = T_1+\Delta T"

"T_2 = 20.0+5.0 = 25.0 ^oC"